OpenStudy (anonymous):

Find the vertex of the quadratic function y = x2 + 12x + 40.

OpenStudy (raden):

to get x-vertex, use the formula x = -b/2a here known the values of a = 1 and b = 12 so, x = -12/(2*1) = -12/2 = -6 now, to get y-vertex just subtitute x=-6 to y(x). it means y(x=-6) = (-6)^2 + 12(-6) + 40 = 36 - 72 + 40 = 4 so, the vertex is at point (-6, 4)